# Extraction of ground-state nuclear deformations from ultra-relativistic heavy-ion collisions: Nuclear structure physics context

### v1

@@ Abstract
The collective-flow-assisted nuclear shape-imaging method in ultra-relativistic heavy-ion collisions has recently been used to characterize nuclear collective states. In this paper, we assess the foundations of the shape-imaging technique employed in these studies. We conclude that, on the whole, the discussion regarding low-energy nuclear physics is confusing and the suggested impact on nuclear structure research is overstated. Conversely, efforts to incorporate existing knowledge on nuclear shapes into analysis pipelines can be beneficial for benchmarking tools and calibrating models used to extract information from ultra-relativistic heavy ion experiments.
@@ I Introduction
The shapes of nuclei colliding in ultra-relativistic heavy-ion collisions (UHIC) influence the geometry of the quark-gluon plasma (QGP) created which, in turn, affects the momentum distribution of the particles produced [1, 2]. The premise of the UHIC studies is rooted in the fact that the resulting anisotropic (hydrodynamic) expansion of the QGP [3] converts the initial spatial asymmetries into momentum anisotropies of the measured particles. A typical method of analyzing this flow consists of characterizing the azimuthal anisotropy by the Fourier coefficients of the flow; the elliptic flow, , is sensitive to the ellipticity of the QGP, while the triangular one, , is sensitive to the triangularity, and so on.
@@
By relating the flow anisotropies to the geometry of the QGP initial state, a shape-imaging method based on UHIC experiments at the Relativistic Heavy Ion Collider (RHIC) and the Large Hadron Collider (LHC) has been suggested as a tool to study the geometrical shapes of colliding nuclei [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. In the present study, we look deeper into the assumptions behind the proposed shape-imaging technique. In particular, we critically assess the claim in Ref. [6] that this approach “not only refines our understanding of the initial stages of the collisions but also turns high-energy nuclear experiments into a precision tool for imaging nuclear structures.”
@@
The paper is organized as follows. Section II discusses the concept of snapshots behind the shape-imaging method, and Sec. III covers the notion of “shape fluctuations.” In Sec. IV we recall the vast body of available knowledge on nuclear shapes from low-energy measurements. The questions related to the uncertainty quantification (UQ) of the shape-imaging method are raised in Sec. V. The conclusions of the paper are given in Sec. VI.
II Symmetries, time scales, and snapshots
Quantum mechanics frequently defies simple intuition. At the same time, simple, intuitive models based on experience often help us to understand the world. These two points should not be confused. Recently, several reports claiming a rigorous determination of nuclear shapes from UHIC experiments have done so to a greater or lesser extent [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19].
@@ II Symmetries, time scales, and snapshots
A fundamental theorem of quantum mechanics states that an eigenstate of a symmetry-conserving Hamiltonian must belong to a representation of the symmetry group. Therefore, the ground states of 238U and 20Ne, which a rotationally-invariant nuclear Hamiltonian governs, belong to the representation, i.e., they are invariant under spatial orientation and reflections. In other words, they are perfectly spherical states in the beam’s reference frame (or perfectly axial in the laboratory frame due to the huge relativistic contraction). Similarly, any ground or excited state with [20, 21] for which the laboratory quadrupole moment is non-zero is covariant under spatial orientation (its rotation mixes the magnetic substates in a specific way). We focus on the case for the following discussion, but note that all arguments apply to any of the others.
@@
A false friend in this understanding is an intuition by which one may picture the deformed 238U nucleus as a rotating and vibrating classical object with corresponding time scales inferred from the energies of the quantum rotational and vibrational states [4]. This is a misleading picture – the ground state of 238U studied in the experiments is a stationary eigenstate, it does not rotate or vibrate, it does not change with time, and, thus, no instantaneous “snapshot” [11] of the shape can be taken, no matter the time scale of the experiment performed.
A proper quantum-mechanical intuition that helps us to understand the shape of the 238U 0+ ground state is a model of a spontaneously symmetry-broken deformed wave function. This is not the wave function that reaches the detector in the UHIC experiment. However, this deformed wave function can be used to build an intuitive model of the 238U 0+ ground state by constructing a linear combination of wave functions differently oriented in space and having different shapes. This model of the 0+ ground state is still time-independent; it does not rotate or vibrate. Most importantly, one must remember that in this model state the wave functions of different orientations and shapes are mixed, not their squares or density distributions, as pictured, e.g., in the tip-tip or body-body configuration schematics of Ref. [5].
@@
So, how then can we measure the shape of a spontaneously symmetry-broken state of 238U or 20Ne? We cannot. The reason is simple: such a stationary state does not exist in nature and, therefore, is unavailable for experimentation. In quantum mechanical modeling, the deformed symmetry-broken state is a wave packet, a coherent superposition of states with different angular momenta. A possible experimental situation where such a state might briefly appear is in a heavy-ion fusion-evaporation reaction, where two ions fuse and rotate together at very high angular momentum. However, immediately after the first de-excitation photon is detected in a -ray detector, the wave packet collapses into a high-angular-momentum eigenstate, which then emits a series of photons and rapidly relaxes to the ground state. Such states are not the ones probed in the UHIC experiments, however.
@@
At this point, we are ready to discuss what is measured in the UHIC experiments. The key point is that one does not measure nuclear moments in these events. Instead, through event reconstruction, the positions of individual nucleons are identified based on measurements of outgoing particles. Multiple identical collisions then provide access to different quantum probability distributions. In short, every UHIC takes a snapshot of the spatial positions of the nucleons, not a snapshot of nuclear rotations or vibrations [11]. In analogy, this can be related to Coulomb explosion experiments, which yield geometrical images of individual molecules [23, 24], see also Ref. [4].
@@
Any measurement of the probability distribution to find one nucleon at any given position in space maps the (exact) one-body density of the nucleus. It is equal to the modulus squared of the (exact) many-body wave function, which depends on the coordinates of particles, integrated over coordinates. The one-body density of any 0+ state is spherical. To determine the shape, one must use the concept of conditional probability, that is, the probability of finding one nucleon at a specific position in space under the condition that another nucleon occupies another position in space. Such a conditional probability is related to the (exact) two-body density (the integral of the exact wave function’s modulus squared over coordinates). It has an axial shape because it must be symmetric with respect to the line connecting the two spatial points in question. In this way, the exact axial shape of a 0+ state can be inferred. Therefore, if the UHIC measurements can extract such conditional probabilities, the axial deformation can be assessed, also in a 0+ state.
@@
Examples of such theoretical analyses, involving two-particle correlations, can be found in, e.g., Refs. [25, 26], where the mean-field, symmetry-breaking, and symmetry-restored solutions are compared with the exact solutions of the problem, with many aspects of the arguments mentioned above clearly and explicitly presented. See also examples of calculations in Refs. [27, 16] of correlated samplings of nucleon positions. In the context of the present discussion, it is useful to note that a local approximation to the two-particle correlation function, based on the density-matrix expansion of the two-body density, is the nucleonic localization function [28], which was originally introduced to visualize bond structures in molecules. This nucleonic localization function was employed in mean-field calculations to study cluster configurations in nuclei [29, 30], cluster structures in fission [31], formation of clusters in both heavy-ion reactions [32] and high-spin states [33].
@@
In the same way, to calculate or measure triaxial deformations, one must consider the (exact) three-body density or measure correlations between positions of three nucleons. This is not what was done in Ref. [12], in which the triaxiality of 129 Xe was inferred based on LHC data.
@@ IV Nuclear deformations: a concept not entirely unknown
Nuclear deformations are parameters characterizing the anisotropy of the nuclear shape. These characteristics are not fundamental nuclear properties, as they are deduced from various measured observables in a model-dependent way [34]. In low-energy nuclear physics, nuclear shapes have been characterized for decades using observables, such as quadrupole moments, obtained from a wide variety of experimental techniques. Since the seminal works by Schüler and Schmidt [35] and Casimir [36, 37], nuclear electric quadrupole moments have been extracted, evaluated, and tabulated for nuclear ground states or excited states with total angular momenta of , from a number of high-precision measurements [38, 39], often quoted with uncertainties of 1% or less for stable nuclei. The nuclear quadrupole moment measures the deviation of the nuclear charge distribution from sphericity in a given nuclear state. In fact, Ref. [38] lists no less than 50 experimental techniques that form the basis for the moments compiled. These range from the hyperfine splitting in mesic atoms [40, 41, 42], to a variety of laser spectroscopic techniques that are even applicable to short-lived rare isotopes [43, 44], to scattering approaches such as electron scattering [45, 46, 47] or heavy-ion-induced Coulomb excitation [48, 49]. For a recent review on the history of nuclear shapes and experimental methods, the reader is referred to the recent review [50].
@@
This body of precision data from low-energy nuclear physics experiments seems largely unknown to the UHIC community and its existence is in direct contradiction with statements such as found in Ref. [12]: “Additionally, for odd-mass and odd-odd nuclei, quantum mechanics entangles the contributions from collective intrinsic deformation and individual single-particle states to the quadrupole moment, such that the former cannot be uniquely determined from a spectroscopic experiment.” One example from Ref. [38] is the moment, b, quoted for the ground state of the odd-mass 197Au nucleus from muonic x-ray spectroscopy, a nucleus that was much discussed by the UHIC community, e.g., in Refs. [13, 4]. Additionally, taking any odd-odd nucleus, here 28Na, from Ref. [38] reveals that moments can be uniquely determined from spectroscopic experiments, e.g., for 28Na from -NMR spectroscopy where a high degree of nuclear polarization was achieved via optical pumping with a collinear laser beam [51]. We note that this was accomplished for a short-lived rare isotope which cannot be accessed by UHIC at RHIC or the LHC.
@@
For most stable and many short-lived nuclei, including for odd-mass and even-even ones, which have a zero ground-state total angular momentum, a vast body of low-energy data exists from Coulomb excitation [48, 49] as well as from electron scattering [52, 53], sometimes even from Mössbauer spectroscopy [54], on the deformation of excited states, including for quadrupole [55, 56] and higher-order deformations such as the octupole (see Refs. [57, 58]) and hexadecapole ones (e.g., [59, 60, 61, 62]). When appropriate in the scientific context, the results from such low-energy measurements use the , , and [63, 48] parameterizations of the deformation as adopted by the UHIC community. Often, the direct use of reduced electromagnetic transitions strengths, , , and proved to be more insightful and scientifically appropriate in the context of the field. The case has yet to be made for a unique nuclear structure insight from UHIC on 238U [4, 14, 64], 20Ne [16, 17], 96Zr and 96Ru [15], 129Xe [12], 197Au [13], and 150Nd [18], given the substantial and quantum-mechanically correctly interpreted body of low-energy data on nuclear deformation [55, 38, 56].
@@
Theoretically, nuclear deformations can be obtained from calculated one-body nuclear densities of symmetry-broken states, including charge density, proton density, or neutron density. A substantial body of literature exists on local and global predictions of nuclear shape deformations, primarily obtained with various flavors of mean-field theory and its extensions. Global surveys can be found in, e.g., Refs. [65, 66, 67, 68, 69, 70, 71, 72, 73] and the theoretical databases MassExplorer [74] and BrusLib [75]. As discussed in Ref. [34], there is a weak model dependence of nuclear deformations of well-deformed nuclei as they reflect the geometries of the valence Hartree-Fock (Kohn-Sham, canonical, Nilsson) orbits. While ab-initio computations have been advocated in several UHIC imaging papers [6, 16], the current state of the ab initio description of deformations is either not satisfactory [76] or in its infancy [77], and more work is needed to make it globally competitive with well-calibrated mean-field-based methods.
@@ V Uncertainty quantification
Uncertainty quantification is at the heart of any pipeline connecting models that span disparate scales. By quantifying the models’ discrepancies and deficiencies, one can construct a rigorous statistical framework that provides a means to transfer information both forwards and backwards within this multiscale (or multifidelity) modeling framework. In the context of the current work, the forward direction involves propagating uncertainties from quantified nuclear models through to final-state observables in UHIC. The natural next step is then to have information flowing in the other direction in the pipeline. This provides an opportunity to solve the inverse problem and perform statistical inference on shapes and other correlated observables accessible within the chosen nuclear structure model. Both steps represent exciting developments and are excellent examples of the confluence of advances in theoretical modeling, experimental analyses, and computational statistics and data science in modern nuclear physics.
@@
Centrality, for instance, in heavy-ion collisions is typically determined by measuring the number of charged particles produced in the collision and then comparing this number to a Monte Carlo Glauber calculation or another similar framework [85, 86, 87, 88]. This approach, while widely used, has certain limitations. The accuracy of centrality determination relies on the validity of the chosen model, which has inherent assumptions about the nucleon-nucleon interactions and the density distribution of nucleons within the nucleus. Additionally, the efficiency of detecting charged particles and the methods used to distinguish between particles originating from the collision and those from secondary decays can introduce uncertainties. The subsequent hydrodynamic evolution of the QGP is another example as it relies on equations of state and transport coefficients that are not fully constrained by experimental data, and even then, there are multiple competing models to propagate the plasma after formation [89]. Finally, the process of hadronization, whereby the QGP converts back into hadrons, introduces further uncertainties related to the fragmentation, clusterization, and subsequent decay of these particles [90, 91].
The complexity and non-uniqueness of this modeling framework raise concerns regarding the robustness and reproducibility of results drawn from the method. Different research groups, employing subtly different implementations of the same underlying stages, can arrive at different conclusions regarding the extracted deformations of the colliding system. These discrepancies can arise from variations in the specific data sets considered in the fit, the choice of initial conditions, the details of the hydrodynamic evolution, the treatment of hadronization, and other factors.
@@
For the reverse direction, the core question is “are the constraints and resulting uncertainties sufficient to distinguish between models used to predict the initial state?”. If the discussed pipeline is rather coarsely constraining, then one should only expect differences to arise from large-scale, bulk changes in the nucleus. For models meeting the reasonableness criteria above, these bulk properties are likely to be very similar. In particular, employing ab-initio nuclear structure models for inference on nuclear properties of any sort requires quantified Hamiltonians and detailed model discrepancy studies to validate those models on nuclear deformations and to render any conclusions statistically significant. While this discussion has focused on nuclear shapes, this is a core point for the extraction of any properties of nuclei. For example, the investigation in Ref. [18] suggests UHIC experiments as a novel probe of matrix elements central to the search for neutrinoless double-beta decay, although without considering the constraining potential of those UHIC observables in the context of other available nuclear structure observables (see, e.g., Ref . [92] for a multitude of factors impacting the matrix element for neutrinoless double-beta decay). More fundamentally, one needs to demonstrate the suitability of the structure model itself for such studies by showing that, when considering data from other sources, the results are not in conflict with well-known nuclear properties.

### v3

@@ Abstract
The collective-flow-assisted nuclear shape-imaging method in ultra-relativistic heavy-ion collisions has recently been used to characterize nuclear collective states. In this paper, we assess the foundations of the shape-imaging technique employed in these studies. We argue that some current UHIC nuclear imaging techniques neglect fundamental aspects of spontaneous symmetry-breaking and symmetry-restoration in colliding ions and incorrectly infer one-body multipole moments from studies of nucleonic correlations. Therefore, the impact of this approach on nuclear structure research has been overstated. Conversely, efforts to incorporate existing knowledge on nuclear shapes into analysis pipelines can be beneficial for benchmarking tools and calibrating models used to extract information from ultra-relativistic heavy-ion experiments.
@@ I Introduction
The shapes of nuclei colliding in ultra-relativistic heavy-ion collisions (UHIC) influence the geometry of the quark-gluon plasma (QGP) created, which, in turn, affects the momentum distribution of the particles produced [1, 2, 3]. The premise of the UHIC studies is rooted in the fact that the resulting anisotropic (hydrodynamic) expansion of the QGP [4] converts the initial spatial asymmetries into momentum anisotropies of the measured particles. A typical method of analyzing this flow consists in characterizing the azimuthal anisotropy by the Fourier coefficients of the flow [5]; the elliptic flow, , is sensitive to the ellipticity of the QGP, while the triangular one, , is sensitive to the triangularity, and so on.
@@
By relating the flow anisotropies to the geometry of the QGP initial state, a shape-imaging method based on UHIC experiments at the Relativistic Heavy Ion Collider (RHIC) and the Large Hadron Collider (LHC) has been suggested as a tool to study the geometrical shapes of the colliding nuclei [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. In the present study, we examine the assumptions underlying the proposed shape-imaging technique in greater detail. In particular, we critically assess the claim in Ref. [17] that this approach “not only refines our understanding of the initial stages of the collisions but also turns high-energy nuclear experiments into a precision tool for imaging nuclear structures.”
@@
The paper is organized as follows. Section II discusses essential concepts of symmetry breaking and restoration in the context of the two-body densities and Sec. III covers the notion of “shape fluctuations.” In Sec. IV we recall the vast body of available knowledge on nuclear shapes from low-energy measurements. The questions related to the uncertainty quantification (UQ) of the shape-imaging method are raised in Sec. V. The conclusions of the paper are given in Sec. VI.
II Symmetries: breaking and restoration
@@
Let us first examine the symmetry aspects of self-bound many-body systems such as atomic nuclei. A fundamental theorem of quantum mechanics states that an eigenstate of a symmetry-conserving Hamiltonian must belong to a representation of the symmetry group. Therefore, the ground states of 238U and 20Ne, which a rotationally-invariant nuclear Hamiltonian governs, belong to the representation, i.e., they are invariant under spatial orientation and reflections. In other words, they are perfectly spherical states in the beam’s reference frame (or perfectly axial in the laboratory frame due to the huge relativistic contraction). Similarly, any ground or excited state with [24, 25] for which the laboratory quadrupole moment is non-zero is covariant under spatial orientation (its rotation mixes the magnetic substates in a specific way). We focus on the case for the following discussion, but note that all arguments apply to any of the others.
@@
A false friend in this understanding is an intuition that allows one to picture the deformed 238U nucleus as a rotating and vibrating classical object, with corresponding time scales inferred from the energies of the quantum rotational and vibrational states [10, 12, 16, 17, 18]. This is a misleading picture – the ground state of 238U studied in the experiments is a stationary eigenstate; it does not rotate or vibrate, it does not change with time, and, thus, no instantaneous snapshot of the shape can be taken, no matter the time scale of the experiment performed.
A proper quantum-mechanical intuition that helps to understand the shape of the 238U 0+ ground state is a model of a spontaneously symmetry-broken deformed wave function. This is not the wave function that reaches the detector in the UHIC experiment. However, this deformed wave function can be used to build an intuitive model of the 238U 0+ ground state by constructing a linear combination of wave functions differently oriented in space and having different shapes, that is, by restoring the symmetry [27]. This model of the 0+ ground state is still time-independent; it does not rotate or vibrate. Most importantly, one must remember that in this model state, the wave functions of different orientations and shapes are mixed, not their squares or density distributions, as shown in diagrams like the tip-tip or body-body configurations of Ref. [18]. Recently, this aspect has been further explored quantitatively in Ref. [28].
@@
So, how then can one measure the shape of a spontaneously symmetry-broken state of 238U or 20Ne? One cannot. The reason is simple: such a stationary state does not exist in nature and, therefore, is unavailable for experimentation. In quantum mechanical modeling, the deformed symmetry-broken state is a wave packet, a coherent superposition of states with different angular momenta. A possible experimental situation where such a state might briefly appear is in a heavy-ion fusion-evaporation reaction, where two ions fuse and rotate together at very high angular momentum. However, immediately after the first de-excitation photon is detected in a -ray detector, the wave packet collapses into a high-angular-momentum eigenstate, which then emits a series of photons and rapidly relaxes to the ground state. In conclusion, the broken-symmetry states are not the ones examined in the UHIC experiments.
@@
At this point, we are ready to discuss what is measured in the UHIC experiments. The key point is that one does not measure one-body nuclear moments in these events. Instead, through event reconstruction, the positions of individual nucleons are identified based on measurements of outgoing particles. Multiple identical collisions then provide access to different quantum probability distributions. In short, every UHIC takes a snapshot of the spatial positions of the nucleons, not a snapshot of nuclear rotations or vibrations. In analogy, this can be related to Coulomb explosion experiments, which yield geometrical images of individual molecules [29, 30].
@@
Any measurement of the probability distribution to find one nucleon at any given position in space maps the (exact) one-body density of the nucleus. It is equal to the modulus squared of the (exact) many-body wave function, which depends on the coordinates of particles, integrated over coordinates. The one-body density of any 0+ state is spherical. To determine the shape, one must use the concept of conditional probability [31], that is, the probability of finding one nucleon at a specific position in space under the condition that another nucleon occupies another position in space. Such a conditional probability is related to the (exact) two-body density (the integral of the exact wave function’s modulus squared over coordinates). It has an axial shape because it must be symmetric to rotations about the line defining the position of the first point in question, and it can provide the density map in the plane spanned by the positions of both points. In this way, the axial shape of a 0+ state can, in principle, be determined. Therefore, if the UHIC measurements can extract such conditional probabilities, the axial deformation can also be assessed in a 0+ state. However, the one-body densities, as used in Refs. [13, 14, 15], are insufficient for that purpose.
Examples of earlier theoretical analyses, involving two-particle correlations, can be found in, e.g., Refs. [31, 27, 32], where the mean-field, symmetry-breaking, and symmetry-restored solutions are compared with the exact solutions of the problem, with many aspects of the arguments mentioned above explicitly presented. It is only recently that the nuclear UHIC modelers realized the importance of the many-body nuclear density for determining nucleons’ position distributions Refs. [33, 34, 35], and practical calculations of correlated samplings of nucleon positions were carried out in only one case study [19].
The two-particle densities from UHIC imaging must be compared to those from modeled wave functions. However, this method for experimentally determining nuclear multipole moments is indirect because the values heavily depend on the model, and the impact of this dependence on uncertainty estimates remains unknown. As a result, they are not competitive with other methods that have been used so far, see Sec. IV.
Additionally, the quantum mechanical spectroscopic quadrupole moment derived from the multipole expansion of the two-body density and the one derived from the one-body density are different observables. Notably, the first is measured in UHIC experiments, while the second is studied in low-energy experiments. Specifically, the first is accessible for states, whereas the second is not. Both may provide complementary but not necessarily identical information about nuclear ground states. This fact seems to have been overlooked in all existing experimental and theoretical studies of the problem. Theoretically, for states, both moments can be calculated and compared using symmetry-conserving ab initio or symmetry-restored mean-field methods. The analysis along these lines is in progress [36].
@@
In the context of the present discussion, it is helpful to note that a local approximation to the two-particle correlation function, based on the density-matrix expansion of the two-body density, is the nucleonic localization function [37]. This function was initially introduced to visualize bond structures in molecules. The analogous nucleonic localization function was employed in mean-field calculations to study cluster configurations in nuclei [38, 39], cluster structures in fission [40], and formation of clusters in both heavy-ion reactions [41] and high-spin states [42].
@@
In the same way, to calculate or measure triaxial deformations, one must consider the (exact) three-body density or measure correlations between positions of three nucleons [33]. This differs from what was done in Ref. [11, 13], where nuclear triaxiality was inferred based on UHIC data.
@@ III Static shapes and dynamic shape fluctuations
The effects describing the dispersion of static observables in the ground states of nuclei are well understood in nuclear structure. They are adequately accounted for by symmetry restoration after mixing the symmetry-broken states [43, 44]. Although they often go by the name of the “vibrational” corrections or “fluctuations” (quadrupole, octupole, pairing, etc.), they model the stationary nuclear states. Such an approach to modeling UHIC results is in its infancy, see, e.g., [13, 20, 21]. Still, it is a proper avenue to take, provided it is used to determine two-body densities and not only the multipole moments.
@@ IV Nuclear deformations: a concept not entirely unknown
Nuclear deformations are parameters characterizing the anisotropy of the nuclear shape. These characteristics are not fundamental nuclear properties, as they are deduced from various measured observables in a model-dependent way [45]. In low-energy nuclear physics, nuclear shapes have been characterized for decades using observables, such as quadrupole moments, obtained from a wide variety of experimental techniques. Since the seminal works by Schüler and Schmidt [46] and Casimir [47, 48], nuclear electric quadrupole moments have been extracted, evaluated, and tabulated for nuclear ground states or excited states with total angular momenta of , from a number of high-precision measurements [49, 50], often quoted with uncertainties of 1% or less for stable nuclei. The nuclear quadrupole moment measures the deviation of the nuclear charge distribution from sphericity in a given nuclear state. In fact, Ref. [49] lists no less than 50 experimental techniques that form the basis for the moments compiled. These range from the hyperfine splitting in mesic atoms [51, 52, 53], to a variety of laser spectroscopic techniques that are even applicable to short-lived rare isotopes [54, 55], to scattering approaches such as electron scattering [56, 57, 58] or heavy-ion-induced Coulomb excitation [59, 60]. For a recent review on the history of nuclear shapes and experimental methods, the reader is referred to Ref. [61].
@@
This body of precision data from low-energy nuclear physics experiments seems to be overlooked by the UHIC community. Ref. [13], for instance, questions the interpretation of spectroscopic experiments for odd- and odd-odd nuclei even though such systems have yielded precise moments which ultimately quantify the deviation from sphericity. One example from Ref. [49] is the moment, b, quoted for the ground state of the odd-mass 197Au nucleus from muonic x-ray spectroscopy, a nucleus that was much discussed by the UHIC community, e.g., in Refs. [14, 16]. Additionally, taking any odd-odd nucleus, here 28Na, from Ref. [49] reveals that moments can be uniquely determined from spectroscopic experiments, e.g., for 28Na from -NMR spectroscopy where a high degree of nuclear polarization was achieved via optical pumping with a collinear laser beam [62]. We note that this was accomplished for a short-lived rare isotope which cannot be accessed by UHIC at RHIC or the LHC.
@@
For most stable and many short-lived nuclei, including even-even nuclei, which have a zero ground-state total angular momentum, and odd-mass systems, a vast body of low-energy data exists from Coulomb excitation [59, 60] as well as from electron scattering [63, 64], sometimes even from Mössbauer spectroscopy [65], on the deformation of excited states, including for quadrupole [66, 67] and higher-order deformations such as the octupole (see Refs. [68, 69]) and hexadecapole ones (e.g., [70, 71, 72, 73]). When appropriate in the scientific context, the results from such low-energy measurements use the , , and [74, 59] parameterizations of the deformation as adopted by the UHIC community. Often, the direct use of reduced electromagnetic transitions strengths, , , and proved to be more insightful and scientifically appropriate in the context of the field. The case has yet to be made for a unique nuclear structure insight from UHIC on 238U [16, 15, 75], 20Ne [19, 20], 96Zr and 96Ru [12], 129Xe [13], 197Au [14], and 150Nd [21], given the substantial and consistently evaluated body of low-energy data on nuclear deformation [66, 49, 67].
@@
Theoretically, nuclear deformations can be obtained from calculated one-body nuclear densities of symmetry-broken states, including charge density, proton density, or neutron density. A substantial body of literature exists on local and global predictions of nuclear shape deformations, primarily obtained with various flavors of mean-field theory and its extensions. Global surveys can be found in, e.g., Refs. [76, 77, 78, 79, 80, 81, 82, 83, 84] and the theoretical databases MassExplorer [85] and BrusLib [86]. As discussed in Ref. [45], there is a weak model dependence of nuclear deformations of well-deformed nuclei as they reflect the geometries of the valence Hartree-Fock (Kohn-Sham, canonical, Nilsson) orbits.
@@ V Uncertainty quantification
Uncertainty quantification is at the heart of any pipeline connecting models that span disparate scales. By quantifying the models’ discrepancies and deficiencies, one can construct a rigorous statistical framework that provides a means to transfer information both forwards and backwards within this multiscale (or multifidelity) modeling framework. In the context of the current work, the forward direction involves propagating uncertainties from quantified nuclear models through to final-state observables in UHIC. The natural next step is then to have information flowing in the other direction in the pipeline. This provides an opportunity to solve the inverse problem and perform statistical inference on shapes and other correlated observables accessible within the chosen nuclear structure model. Both steps represent exciting developments and are examples of the confluence of advances in theoretical modeling, experimental analyses, and computational statistics and data science in modern nuclear physics.
@@
Centrality, for instance, in heavy-ion collisions is typically determined by measuring the number of charged particles produced in the collision and then comparing this number to a Monte Carlo Glauber calculation or another similar framework [94, 95, 96, 97]. This approach, while widely used, has certain limitations. The accuracy of centrality determination relies on the validity of the chosen model, which has inherent assumptions about the nucleon-nucleon interactions and the density distribution of nucleons within the nucleus. Additionally, the efficiency of detecting charged particles and the methods used to distinguish between particles originating from the collision and those from secondary decays can introduce uncertainties. The subsequent hydrodynamic evolution of the QGP is another example, as it relies on equations of state and transport coefficients that are not fully constrained by experimental data, and even then, there are multiple competing models to propagate the QGP after formation [98]. In addition to this hydrodynamical flow, there are also nonflow correlations that must be corrected for and that add further model-based uncertainties to the analysis pipeline [99]. The recent debate [100, 101] underscores the importance of these nonflow contributions, specifically in the context of nuclear shape extraction. Finally, the process of hadronization, whereby the QGP converts back into hadrons, introduces further uncertainties related to the fragmentation, clusterization, and subsequent decay of these particles [102, 103].
The complexity and non-uniqueness of this modeling framework raise concerns regarding the robustness and reproducibility of results drawn from the method. Different research groups, employing subtly different implementations of the same underlying stages, could potentially arrive at different conclusions regarding the extracted deformations of the colliding system. These discrepancies can arise from variations in the specific data sets considered in the fit, the choice of initial conditions, the details of the hydrodynamic evolution, the treatment of hadronization, and other factors. While this model dependence has not been investigated in detail for shape extraction, model discrepancy studies have been performed for UHIC-related quantities in Ref. [104], and the impact on the Bayesian inference was found to be significant.
@@
For the reverse direction, the core question is “are the constraints and resulting uncertainties sufficient to distinguish between models used to predict the initial state?”. If the discussed pipeline is rather coarsely constraining, then one should only expect differences to arise from large-scale, bulk changes in the nucleus. For models meeting the reasonableness criteria above, these bulk properties are likely to be very similar. Furthermore, employing any nuclear structure models for inference on nuclear properties of any sort requires quantified Hamiltonians and detailed model discrepancy studies to validate those models on nuclear deformations and to render any conclusions statistically significant. This is equally true for any model of nuclear structure, including ab initio models. While this discussion has focused on nuclear shapes, this is a core point for the extraction of any properties of nuclei. For example, the investigation in Ref. [21] suggests UHIC experiments as a novel probe of matrix elements central to the search for neutrinoless double-beta decay, although without considering the constraining potential of those UHIC observables in the context of other available nuclear structure observables (see, e.g., Ref . [105] for a multitude of factors impacting matrix elements for neutrinoless double-beta decay). More fundamentally, one needs to demonstrate the suitability of the structure model itself for such studies by showing that, when considering data from other sources, the results are not in conflict with well-known nuclear properties.